At class meeting, Jeff, Chad and Lane ate a total of 254 marshmallows in the contest. Jeff won by eating 7 more than Chad, who ate 5 more than Lane. How many marshmallows did Jeff eat?
Answer by ewatrrr(24785)


Hi
x Lane's amount
Jeff + Chad + Lane = 254
(x+12)+(x+5)+ x = 254
3x = 237
x = 79, Lane ate. Jeff 91
And
91 + 84 + 79 = 254 checks!